OpenStudy (anonymous):

how much water is formed by mixing 100g h2 and 100g o2????

OpenStudy (chmvijay):

can you first write the balanced reaction between the H2 and O2 to form H20

OpenStudy (anonymous):

yess 2h2+o2=2h2o

OpenStudy (anonymous):

then what is the next step

OpenStudy (chmvijay):

here you have to find out of these two reactant which one is in excess :)

OpenStudy (anonymous):

kk thwn how will we do it

OpenStudy (anonymous):

i thnk h20 i excess

OpenStudy (chmvijay):

nooo H2O is product i am asking from the reactant

OpenStudy (chmvijay):

mean 4 gram of H2 react with 32 gram ofO2 to produce 36 g of water now can you tell which is present in excess reactant if you consider both 100 g

OpenStudy (anonymous):

then o2

OpenStudy (chmvijay):

yaaa now 32 gram of O2 will form 36 gram of H2O then100 gram of O2 will form =36*100 /32 = can u calculate

OpenStudy (chmvijay):

hey did u calculate that

OpenStudy (anonymous):

112 hai answer

OpenStudy (anonymous):

yaaa got it thnkss

OpenStudy (anonymous):

answer is 112.5